A balanced delta connected load is supplied from a 220V, 60Hz source thru a line with impedance of 0.5 + j0.8 ohm. If the power consumed by the load is 3 KW at 0.8 power factor leading. Determine the line currents and phase currents in phasor form. CBA Sequence
1
Expert's answer
2020-10-08T15:13:04-0400

Since it is necessary to find currents in the phase form, then to begin with we find the phase current through the ratio "I_{phasor}=\frac U z" , where "U" is the voltage, "z" is the impedance, the voltage in the phase form "U = 220\times e ^ {j\times 0}" ;

impedance in phase form "z = \sqrt {0.89 }\times e ^ {j\times atan (\frac 8 5)}" therefore

"I_{phasor}=\frac {220\times e^{j\times 0}} {\sqrt {0.89}\times e^{j\times atan(\frac 8 5)}}=123.6-j \times 197.75" A;

Since the phase current and the linear are related by the relation "I_{line}=I_{phasor}\times \sqrt {3}" , then

"I_{line}=(123.6-j \times 197.75) \times \sqrt{3}" "=214.08-j \times 342.51" A.